# A Test of Time: Predicting the Sustainable Success of Online Collaboration in Wikipedia

### v1

@@ Abstract.
The Internet has significantly expanded the potential for global collaboration, allowing millions of users to contribute to collective projects like Wikipedia. While prior work has assessed the success of online collaborations, most approaches are time-agnostic, evaluating success without considering its longevity. Research on the factors that ensure the long-term preservation of high-quality standards in online collaboration is scarce. In this study, we address this gap. We propose a novel metric, ‘Sustainable Success,’ which measures the ability of collaborative efforts to maintain their quality over time. Using Wikipedia as a case study, we introduce the SustainPedia dataset, which compiles data from over 40K Wikipedia articles, including each article’s sustainable success label and more than 300 explanatory features such as edit history, user experience, and team composition. Using this dataset, we develop machine learning models to predict the sustainable success of Wikipedia articles. Our best-performing model achieves a high AU-ROC score of 0.88 on average. Our analysis reveals important insights. For example, we find that the longer an article takes to be recognized as high-quality, the more likely it is to maintain that status over time (i.e., be sustainable). Additionally, user experience emerged as the most critical predictor of sustainability. Our analysis provides insights into broader collective actions beyond Wikipedia (e.g., online activism, crowdsourced open-source software), where the same social dynamics that drive success on Wikipedia might play a role. We make all data and code used for this study publicly available for further research.
Crowd collaboration, Computational Social Science, Online Communities, Wikipedia
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Preprint; arXiv; October 2024isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing User studiesccs: Information systems Wikisccs: Information systems Social networksccs: Applied computing Sociologyccs: Social and professional topics Cultural characteristics
1. Introduction
Collaboration plays a crucial role in shaping societies (Sachs et al., 2004; Salminen, 2012; Ostrom, 1990). Online platforms like Wikipedia and Github enable crowd collaboration, allowing millions of users to collectively create content (Ebersbach et al., 2008; Godwin-Jones, 2003; Lima et al., 2014; Dabbish et al., 2012; Piccardi et al., 2021). Yet, not all collaborations are successful. Studies focused on how to measure, predict, and evaluate the success of crowd collaboration has gained considerable attention in recent years. Typically, the focus is driven by two key motivations: (i) Better predicting the success of online collaborations (Moás and Lopes, 2023); and (ii) Gaining insights and a clearer understanding of the factors that contribute to success (Platt and Romero, 2018; Israeli and Tsur, 2024; Luther et al., 2010). However, most proposed approaches focus on success at a single point in time, overlooking the importance of sustained success over the long term—i.e., what factors contribute to an initially successful collaboration falling apart? To address this gap, we introduce a novel metric called ’Sustainable Success,’ which assesses the ability of online collaborations to maintain high-quality standards over time.
@@ 1. Introduction
Using our sustainable success measure, this study asks the following two research questions:
(RQ1) Can the sustainable success of an online collaboration be accurately predicted?; and (RQ2) What factors contribute to the sustainable success of an online collaboration?
@@
To answer these questions, we measure collaborations on over a decade of behavior in Wikipedia. Through collaborative efforts, Wikipedia’s content stands as one of the largest free knowledge repositories in the world (similarweb, 2024; Ren et al., 2023). Wikipedia has a rich collaborative environment, with various means of user interaction, including voting, discussions, and communication. Crucially, Wikipedia editors also directly evaluate its content in a structured way (Hu et al., 2007; Stvilia et al., 2005; Warncke-Wang et al., 2013; Zhang et al., 2017), providing a reliable basis for measuring sustainable success. The public accessibility of Wikipedia’s data allows us to comprehensively explore the factors that contribute to the sustainability of collaborative efforts, thereby answering our RQs in a thorough and reliable way.
@@
In our study, we use Wikipedia’s two most trustworthy content assessment levels: Featured Article ($FA$) and Good Article ($GA$). The two tags signify high-quality articles that meet specific editorial standards (Wikipeida, 2024c, e). $FA$ represents the pinnacle of quality, having undergone rigorous peer review and meeting the strictest criteria, such as comprehensive coverage, neutrality, and well-researched content. $GA$s, while not as rigorous as $FA$s, also meet substantial quality standards, including being well-written, factually accurate, and verifiable. As of 10/2024, just 0.1% and 0.5% of English Wikipedia articles are tagged as $FA$ and $GA$, respectively.
@@
The community’s judgment of quality is not permanent, however. The $FA$ and $GA$ statuses can both be removed if the article no longer meets the community’s quality standards, effectively demoting the article. Figure 1 illustrates this promotion and demotion process. We leverage this ongoing quality assessment to model the sustainable success of Wikipedia articles. We frame the problem as a classification task, categorizing articles as either demoted (unsustainable) or not, in order to develop machine learning models for predicting sustainable success. To understand the process behind sustained success, our models use different categories of factors related to creating and maintaining the article, including team composition, discussion patterns, and the experience of the article’s contributors. To summarise, our main three contributions are:
  1. (1)
@@
Figure 1. Article life-cycle example. (Top) Major events of the Wikipedia article about the edible fruit Durian. (Bottom) A screenshot from the talk page of the Durian article, including both its promotions and demotion. This article is tagged as unsustainable in SustainPedia due to its demotion in 2022.
2. Online Collaboration and Success
Numerous researchers investigate collaborative behaviors on social platforms like Reddit and X (formerly Twitter) (Honey and Herring, 2009; Weninger et al., 2013; Stoddard, 2015; Panek et al., 2018; Mensah et al., 2020). Additionally, some focus on identifying these collaborative patterns within specialized crowd-sourced platforms, including Q&A websites, open-source initiatives, and Wikipedia (Wang et al., 2013; Ransbotham and Kane, 2011; Dabbish et al., 2012; Keegan et al., 2012; Thung et al., 2013; Zhang et al., 2017).
A significant area of focus in this field is quantifying and predicting the success of online activities (Tan, 2018; Cunha et al., 2019; Israeli and Tsur, 2024), particularly in relation to crowd collaboration (Das et al., 2022, 2024). Many of such efforts focus on understanding current algorithms used for quality assessment in Wikipedia (Ren et al., 2023). These success metrics, among other things, help identify the elements that contribute to successful online collaboration such as edit behavior and group dynamics.
In the past decade, numerous studies have demonstrated how factors such as edit behavior (Liu and Ram, 2011; Ruprechter et al., 2020), network structure (Raman et al., 2020; Brandes et al., 2009; Lerner and Lomi, 2019; de La Robertie et al., 2015), regulation (Ruprechter et al., 2023), and textual and visual content (Bassani and Viviani, 2019; Shen et al., 2017, 2019; Guda et al., 2020) are associated with content assessment in Wikipedia. Some of those suggest advanced deep learning models that integrate multiple data sources using a multimodal architecture.
@@ 3. Problem Definition
3.1. Wikipedia’s Quality Assessment System
Articles are assigned quality scores in Wikipedia in two ways. First, Wikipedia maintains two quality assessment classes where an article can be tagged as “Featured Article” ($FA$) or “Good Article” ($GA$). $FA$ and $GA$ determinations are made using a Wikipedia broad consensus. $FA$s are “the best articles Wikipedia has to offer” (Wikipeida, 2024c) while $GA$s meet “a core set of editorial standards” (Wikipeida, 2024e) but are not featured article quality. Second, quality grades (e.g., B, C, Start) are assigned by WikiProjects, which are groups of thematically related articles, such as the ‘Plants’ Wikiproject. Most articles are associated with at least one WikiProject who each assesses an article’s quality separately using the same grading scale. However, each WikiProject can have different quality standards, grading systems, responsiveness to changes in an article, etc. It is common for an article to have different quality grades across WikiProjects (openZIM, 2024).
In contrast, $FA$ and $GA$ quality assessments are judged Wikipedia-wide using a systematic peer-review evaluation, which excludes the editors of the evaluated article, making them more reliable than WikiProjects scores. Thus, in this study, we only use $FA$ and $GA$ quality assessment levels. Currently, there are 6,583 and 40,191 $FA$s (Wikipeida, 2024c) and $GA$s (Wikipeida, 2024e) respectively.
3.2. Promotions and Demotion
@@
The demotion of an article, the same as promotion, is determined through a consensus derived through discussion at a dedicated Wikipedia discussion page. The reevaluation of an article’s quality is done through an ‘article review’ process, which usually takes a few weeks (Wikipeida, 2024b, f). During this time, the members of the relevant community are allowed to comment on the required improvements and directly edit the article to improve it. Appendix Figure 6 illustrates how a review process is announced. A bold comment appears on the ‘Talk Page’ of the relevant article, asking the Wikipedia community to comment about the issue and motivating Wikipedians to edit the article and improve its content.
@@ 3.3. Binary Classification
(1) $\mathbbm{1}_{unsustain}^{a}\triangleq\begin{cases}1&\text{$a$ was promoted andlater demoted}\\ 0&\text{$a$ was promoted and never demoted}\end{cases}$
@@
4. Data
Each Wikipedia article maintains its content page (the actual article) and its talk page. In this research, we make use of both pages. While the content page and its meta-data contain essential information on how the article’s content is built and by whom, the latter contains discussions between the community members. We collect the data of all relevant articles using the Wikipedia dump files, which are monthly released by the WikiMedia Foundation (Wikimedia, 2024).
@@ 4. Data
4.1. SustainPedia Population
Table 1. SustainPedia dataset statistics. $\mathbbm{1}_{unsustain}$ is the classification target feature. ‘Prom. Time’ is $t_{prom}-t_{birth}$, measured in days. Comments are taken from discussions on talk pages. The mean and standard deviation are calculated over articles. Blanks are irrelevant.
@@ 4.1. SustainPedia Population
To construct the SustainPedia population, we look for articles that are currently or have been in the $FA$ or $GA$ status, including those that have been demoted. Wikipedia maintains lists of existing $FA$s and $GA$s (Wikipeida, 2024c, e). In addition, there are Wikipedia pages that list demoted $FA$s, named ‘Former Featured Articles,’ and demoted $GA$s, named ‘Delisted Good Articles’ (Wikipeida, 2024a, d). We use these four lists to construct the set of articles in SustainPedia.
4.2. Target Feature Construction
@@ 4.3. Explanatory Features
This feature set, named $EH$, captures the editing dynamics of the article, which have been shown to be associated with performance in Wikipedia (Ruprechter et al., 2020; Kittur and Kraut, 2008). We extract and calculate $EH$ from the article’s revisions history in the [$t_{birth}^{a}$, $t_{prom}^{a}$] timespan. It includes, number-of-editors ($EH_{1}$), number-of-revisions ($EH_{2}$), time-to-promotion ($EH_{3}$), and reverted-revisions-percentage ($EH_{4}$) which is the number of reverts out of $EH_{2}$. We also create normalized features for the first two metrics to account for edit activity and time: $EH_{1}/EH_{2}$ and $EH_{2}/EH_{3}$, respectively. The Was-a-good-article ($EH_{5}$) binary feature indicates whether the article was first promoted to $GA$ before a second promotion to $FA$ (see example in Figure 1). It is only relevant for $FA$s.
@@
Numerous scholars show the effect of network structure among collaborators and success (Hinds and Lee, 2008; Romero et al., 2015, 2016). In Wikipedia, a few studies focus on the network structures and their impact on Wikipedia articles (Brandes et al., 2009; Jurgens and Lu, 2012; Lerner and Lomi, 2019; Qin et al., 2015). We refer to network features as the ones extracted from the edit-network $G_{a}=(V_{a},E_{a})$ that we construct for each $a\in A$. $G_{a}$ is based on the edit interactions among collaborators. $V_{a}$ (nodes) and $E_{a}$ (edges) represent Wikipedians and their interactions, respectively. [^2] Anonymous editors are excluded. Inspired by Platt and Romero (2018) and Lerner and Lomi (2019), we build $G_{a}$ according to the sequential revision history. A directed edge $e\in E_{a}$ is assigned between ${(v_{1},v_{2})}\in V_{a}$ if $v_{2}$ revised article $a$, directly after a revision by $v_{1}$.
@@
Existing scholars highlight the relation between Wikipedia topics to their edit frequency and edit wars (Rad and Barbosa, 2012; Wilson and Likens, 2015). We hypothesize that incorporating topical features will improve the model’s prediction. These features not only enhance the model’s accuracy but also decrease the likelihood that the associations we find between other features and sustainability are not confounded by the topic of the article.
As we note in Section 3, most Wikipedia articles are associated with multiple WikiProjects. Each WikiProject rates the importance of the article to the WikiProject as ‘Low,’ ‘Mid,’ ‘High,’ or ‘Top’ (openZIM, 2024). We utilize this information to generate topical features for $a\in A$, using a Likert scale ranging from 0 to 4.333A zero value is assigned when an article is not associated with a WikiProject. A value of four is assigned when an article is associated with a WikiProject as ‘Top-Importance.’ There are over 2,000 English WikiProjects and many of them are very small (Wikipeida, 2024g). We only use the top 250 by number of articles. [^4] Three of the 250 topics were manually removed as they were not purely topical. For example, the ‘Spoken Wikipedia:’ https://tinyurl.com/33h98n2p. 97.7% of the articles in the SustainPedia dataset belong to at least one of these 250 WikiProjects.
@@
The way Wikipedians handle their discussions has been shown to be correlated with various success measures (Maki et al., 2017; Stvilia et al., 2005). We expect to find a significant correlation also with sustainable success. For example, we assume that polite and respectful discussions would be positively associated with sustainable success since these types of discussions may be more likely to incorporate diverse perspectives.
We collected all historical discussions between Wikipedians (i.e., discussers) on the talk page of each article and constructed two types of discussion features. The first type are linguistic features. We use pre-trained deep-learning models to measure those. Specifically, we measure the sentiment (Hartmann et al., 2023), formality (Babakov et al., 2023), politeness (Danescu-Niculescu-Mizil et al., 2013; Wang and Jurgens, 2018), toxicity (Logacheva et al., 2022), and certainty (Pei and Jurgens, 2021) for each comment in a discussion. We then aggregate each, using the average, median, and percentage of high values. [^5] We use previously identified thresholds to define high values using each model To account for uneven contributions across the discussers of talk pages that might bias our measures, we also calculate the mean and median values for each discusser and then aggregate these values across all discussers.
The other type of features are related to the structure of the discussions. We list these 15 features in Appendix Table 3. Among those features are number-of-discussers, mean-responded-comments, and the discussers-gini (Gini, 1912), which is a measure of statistical dispersion commonly used to measure inequality. A Gini index of 0.0 indicates perfect equality where all editors contribute the same amount of content (i.e., comments in the case of discussions), while 1.0 indicates perfect inequality where a single individual generates all content.
@@
The way a team of Wikipedians engages with one another and with the edited article is crucial for the team’s success and resilience (Ruprechter et al., 2020). We hypothesize that this is also correlated with sustainable success. This feature set, named $TC$, includes editors-gini ($TC_{1}$) and ip-based-edits-percentage ($TC_{2}$). $TC_{1}$ corresponds to the Gini inequality coefficient (Gini, 1912) among the Wikipedian editors. It is calculated in the same way as for discussers-gini (see previous paragraph) while taking into account article edit distribution among Wikipedians. $TC_{2}$ represents the fraction of revisions done by anonymous editors.

### v2

@@ Abstract
The Internet has significantly expanded the potential for global collaboration, allowing millions of users to contribute to collective projects like Wikipedia. While some of these collaborative efforts see early success, their long-term success is key to creating lasting impact. Despite its importance, this dimension of success remains largely unexplored. Prior work has assessed the success of online collaborations, however, most approaches are time-agnostic, evaluating success without considering its longevity. Research on the factors that ensure the long-term preservation of high-quality standards in online collaboration is scarce. In this study, we address this gap. We propose a novel metric, ‘Sustainable Success,’ which measures the ability of collaborative efforts to maintain their quality over time. Using Wikipedia as a case study, we introduce the SustainPedia dataset, which compiles data from over 40K Wikipedia articles, including each article’s sustainable success label and more than 300 explanatory features such as edit history, user experience, and team composition. Using this dataset, we develop machine learning models to predict the sustainable success of Wikipedia articles. Our analysis reveals important insights. For example, we find that the longer an article takes to be recognized as high-quality, the more likely it is to maintain that status over time (i.e., be sustainable). Additionally, user experience emerged as the most critical predictor of sustainability. Our analysis provides insights into broader collective actions beyond Wikipedia, such as online activism and crowdsourced open-source software, where the same social dynamics that drive success on Wikipedia might play a role.
1 Introduction
Collaboration plays a crucial role in societal progress, shaping the way individuals and groups share knowledge, solve problems, and innovate Sachs et al. (2004); Salminen (2012); Ostrom (1990); Tschannen-Moran (2001). In the digital era, platforms like Wikipedia and GitHub have revolutionized collaboration by allowing millions of users to collectively create content Ebersbach et al. (2008); Godwin-Jones (2003); Lima et al. (2014); Dabbish et al. (2012); Piccardi et al. (2021). However, while collaboration is rooted in human behavior, and online platforms empower millions to do so easily, the reality remains that not all collaborative efforts lead to success.
Humans have long sought to assess group success across various dimensions and contexts Hackman (2002); Woolley et al. (2010); Pentland (2012); Cunha et al. (2019); Kim (2006); Kraut (2012). In particular, evaluating and predicting the success of collaborative groups is important for optimizing work efficiency and gaining deeper insights into human dynamics.
As a result, there has been growing interest in understanding the factors that determine the success or failure of crowd collaboration. Research in this area is primarily motivated by two key objectives: (i) Better predicting the success of online collaborations Moás and Lopes (2023); and (ii) Uncovering the underlying mechanisms and factors that foster success Platt and Romero (2018); Israeli and Tsur (2024); Luther et al. (2010). Advancing this knowledge is essential not only for improving existing collaborative efforts but also for the management, moderation, and design of future collaborative systems that maximize engagement, efficiency, and impact.
However, most proposed approaches focus on success at a single point in time, overlooking the importance of sustained success over the long term, which necessitates continuous effort and ongoing maintenance by the collaborative team. While the importance of maintenance, repair, and long-term care has recently been recognized in innovation studies Vinsel and Russell (2020), it has yet to be fully explored in the context of collaborative efforts.
Sustained success in collaborative environments is a challenge. Early signals of success do not always guarantee long-term viability. Prior work explored factors that influence the persistence of collaborative efforts, such as governance structures Forte et al. (2009), contributor retention Halfaker et al. (2013); Cascaes Cardoso (2017), and social dynamics within online groups Kittur et al. (2007); Kraut and Fiore (2014). However, none of these studies have explored the capacity of collaborative efforts to maintain a high-quality status over time—i.e., avoid an initially successful collaboration falling apart. To address this gap, we introduce a novel metric called ’Sustainable Success,’ which assesses this aspect of success directly – the ability of collaborative teams to maintain high-quality standards over time.
@@ 1 Introduction
Using the new sustainable success measure, this study asks the following two research questions:
(RQ1) Can the sustainable success of an online collaboration be accurately predicted?; and
(RQ2) What factors contribute to the sustainable success of an online collaboration?
@@
To answer these questions, we measure collaborations on over a decade of behavior in Wikipedia. Through collaborative efforts, Wikipedia’s content stands as one of the largest free knowledge repositories in the world similarweb (2024); Ren et al. (2023). Wikipedia has a rich collaborative environment, with various means of user interaction, including voting, discussions, and communication. Crucially, Wikipedia editors also directly evaluate its content in a structured way Hu et al. (2007); Stvilia et al. (2005); Warncke-Wang et al. (2013); Zhang et al. (2017), providing a reliable basis for measuring sustainable success. The public accessibility of Wikipedia’s data allows us to comprehensively explore the factors that contribute to the sustainability of collaborative efforts, thereby answering our RQs in a thorough and reliable way.
@@
In our study, we use Wikipedia’s two most trustworthy content assessment levels: Featured Article ($FA$) and Good Article ($GA$). The two tags signify high-quality articles that meet specific editorial standards Wikipeida (2024b, e). $FA$ represents the pinnacle of quality, having undergone rigorous peer review and meeting the strictest criteria, such as comprehensive coverage, neutrality, and well-researched content.
$GA$s, while not as rigorous as $FA$s, also meet substantial quality standards, including being well-written, factually accurate, and verifiable. As of 02/2025, just 0.1% and 0.49% of English Wikipedia articles are tagged as $FA$ and $GA$, respectively. $FA$s are marked with a bronze star symbol while $GA$s with a green plus. Both tags highlight Wikipedia’s best work, reflecting reliable, in-depth content.
@@
The community’s judgment of quality is not permanent, however. The $FA$ and $GA$ statuses can both be removed if the article no longer meets the community’s quality standards, effectively demoting the article. Figure 1 illustrates this promotion and demotion process. We leverage this ongoing quality assessment to model the sustainable success of Wikipedia articles.
To understand the process behind sustained success, our models use different categories of factors related to creating and maintaining the article. Specifically, we build the following six feature sets: network structure of collaborators, topics that the article is associated with, team composition, discussion patterns, edit history, and the experience of the article’s collaborators.
To summarise, our main three contributions are:
  1. 1.
@@
Figure 1: Article life-cycle example. (Top) Major events of the Wikipedia article about the edible fruit Durian. (Bottom) A screenshot from the talk page of the Durian article, including both its promotions and demotion. This article is tagged as unsustainable in SustainPedia due to its demotion in 2022.
2 Online Collaboration and Success
Numerous researchers investigate collaborative behaviors on social platforms like Reddit and X (formerly Twitter) Honey and Herring (2009); Weninger et al. (2013); Stoddard (2015); Panek et al. (2018); Mensah et al. (2020). Additionally, some focus on identifying these collaborative patterns within specialized crowd-sourced platforms, including Q&A websites, open-source initiatives, and Wikipedia Wang et al. (2013); Ransbotham and Kane (2011); Dabbish et al. (2012); Keegan et al. (2012); Thung et al. (2013); Zhang et al. (2017).
A significant area of focus in this field is quantifying and predicting the success of online activities Tan (2018); Cunha et al. (2019); Israeli and Tsur (2024), particularly in relation to crowd collaboration Das et al. (2022, 2024). Many of such efforts focus on understanding current algorithms used for quality assessment in Wikipedia Ren et al. (2023). These success metrics, among other things, help identify the elements that contribute to successful online collaboration such as edit behavior and group dynamics.
In the past decade, numerous studies have demonstrated how factors such as edit behavior Liu and Ram (2011); Ruprechter et al. (2020), network structure Raman et al. (2020); Brandes et al. (2009); Lerner and Lomi (2019); de La Robertie et al. (2015), regulation Ruprechter et al. (2023), and textual and visual content Bassani and Viviani (2019); Shen et al. (2017, 2019); Guda et al. (2020) are associated with content assessment in Wikipedia. Some of those suggest advanced deep learning models that integrate multiple data sources using a multimodal architecture.
@@ 3 Problem Definition
3.1 Wikipedia’s Quality Assessment System
Articles are assigned quality scores in Wikipedia in two ways. First, Wikipedia maintains two quality assessment classes where an article can be tagged as “Featured Article” ($FA$) or “Good Article” ($GA$). $FA$ and $GA$ determinations are made using a Wikipedia broad consensus. $FA$s are “the best articles Wikipedia has to offer” Wikipeida (2024b) while $GA$s meet “a core set of editorial standards” Wikipeida (2024e) but are not featured article quality. Second, quality grades (e.g., B, C, Start) are assigned by WikiProjects, which are groups of thematically related articles, such as the ‘Plants’ Wikiproject. Most articles are associated with at least one WikiProject who each assesses an article’s quality separately using the same grading scale. However, each WikiProject can have different quality standards, grading systems, responsiveness to changes in an article, etc. It is common for an article to have different quality grades across WikiProjects openZIM (2024).
In contrast, $FA$ and $GA$ quality assessments are judged Wikipedia-wide using a systematic peer-review evaluation, which excludes the editors of the evaluated article, making them more reliable than WikiProjects scores. Thus, in this study, we only use $FA$ and $GA$ quality assessment levels. Currently, there are 6,583 and 40,191 $FA$s Wikipeida (2024b) and $GA$s Wikipeida (2024e) respectively.
3.2 Promotions and Demotion
@@
The demotion of an article, the same as promotion, is determined through a consensus derived through discussion at a dedicated Wikipedia discussion page. The reevaluation of an article’s quality is done through an ‘article review’ process, which usually takes a few weeks Wikipeida (2024c, f). During this time, the members of the relevant community are allowed to comment on the required improvements and directly edit the article to improve it. Appendix Figure 6 illustrates how a review process is announced. A bold comment appears on the ‘Talk Page’ of the relevant article, asking the Wikipedia community to comment about the issue and motivating Wikipedians to edit the article and improve its content.
@@ 3.3 Binary Classification
$\mathbbm{1}_{unsustain}^{a}\triangleq\begin{cases}1&\text{$a$ was promoted andlater demoted}\\ 0&\text{$a$ was promoted and never demoted}\end{cases}$ (1)
@@
4 Data
Each Wikipedia article maintains its content page (the actual article) and its talk page. In this research, we make use of both pages. While the content page and its meta-data contain essential information on how the article’s content is built and by whom, the latter contains discussions between the community members. We collect the data of all relevant articles using the Wikipedia dump files, which are monthly released by the WikiMedia Foundation Wikimedia (2024).
@@ 4 Data
4.1 SustainPedia Population
Table 1: SustainPedia dataset statistics. $\mathbbm{1}_{unsustain}$ is the classification target feature. ‘Prom. Time’ is $t_{prom}-t_{birth}$, measured in days. Comments are taken from discussions on talk pages. The mean and standard deviation are calculated over articles. Blanks are irrelevant.
@@ 4.1 SustainPedia Population
To construct the SustainPedia population, we look for articles that are currently or have been in the $FA$ or $GA$ status, including those that have been demoted. Wikipedia maintains lists of existing $FA$s and $GA$s Wikipeida (2024b, e). In addition, there are Wikipedia pages that list demoted $FA$s, named ‘Former Featured Articles,’ and demoted $GA$s, named ‘Delisted Good Articles’ Wikipeida (2024a, d). We use these four lists to construct the set of articles in SustainPedia.
4.2 Target Feature Construction
@@ 4.3 Explanatory Features
This feature set, named $EH$, captures the editing dynamics of the article, which have been shown to be associated with performance in Wikipedia Ruprechter et al. (2020); Kittur and Kraut (2008). We extract and calculate $EH$ from the article’s revisions history in the [$t_{birth}^{a}$, $t_{prom}^{a}$] timespan. It includes, number-of-editors ($EH_{1}$), number-of-revisions ($EH_{2}$), time-to-promotion ($EH_{3}$), and reverted-revisions-percentage ($EH_{4}$) which is the number of reverts out of $EH_{2}$. We also create normalized features for the first two metrics to account for edit activity and time: $EH_{1}/EH_{2}$ and $EH_{2}/EH_{3}$, respectively. The Was-a-good-article ($EH_{5}$) binary feature indicates whether the article was first promoted to $GA$ before a second promotion to $FA$ (see example in Figure 1). It is only relevant for $FA$s.
@@
Numerous scholars show the effect of network structure among collaborators and success Hinds and Lee (2008); Romero et al. (2015, 2016). In Wikipedia, a few studies focus on the network structures and their impact on Wikipedia articles Brandes et al. (2009); Jurgens and Lu (2012); Lerner and Lomi (2019); Qin et al. (2015). We refer to network features as the ones extracted from the edit-network $G_{a}=(V_{a},E_{a})$ that we construct for each $a\in A$. $G_{a}$ is based on the edit interactions among collaborators. $V_{a}$ (nodes) and $E_{a}$ (edges) represent Wikipedians and their interactions, respectively. [^2] Anonymous editors are excluded. Inspired by Platt and Romero (2018) and Lerner and Lomi (2019), we build $G_{a}$ according to the sequential revision history. A directed edge $e\in E_{a}$ is assigned between ${(v_{1},v_{2})}\in V_{a}$ if $v_{2}$ revised article $a$, directly after a revision by $v_{1}$.
@@
Existing scholars highlight the relation between Wikipedia topics to their edit frequency and edit wars Rad and Barbosa (2012); Wilson and Likens (2015). We hypothesize that incorporating topical features will improve the model’s prediction. These features not only enhance the model’s accuracy but also decrease the likelihood that the associations we find between other features and sustainability are not confounded by the topic of the article.
As we note in Section 3, most Wikipedia articles are associated with multiple WikiProjects. Each WikiProject rates the importance of the article to the WikiProject as ‘Low,’ ‘Mid,’ ‘High,’ or ‘Top’ openZIM (2024). We utilize this information to generate topical features for $a\in A$, using a Likert scale ranging from 0 to 4.333A zero value is assigned when an article is not associated with a WikiProject. A value of four is assigned when an article is associated with a WikiProject as ‘Top-Importance.’ There are over 2,000 English WikiProjects and many of them are very small Wikipeida (2024g). We only use the top 250 by number of articles. [^4] Three of the 250 topics were manually removed as they were not purely topical. For example, the ‘Spoken Wikipedia:’ https://tinyurl.com/33h98n2p. 97.7% of the articles in the SustainPedia dataset belong to at least one of these 250 WikiProjects.
@@
The way Wikipedians handle their discussions has been shown to be correlated with various success measures Maki et al. (2017); Stvilia et al. (2005). We expect to find a significant correlation also with sustainable success. For example, we assume that polite and respectful discussions would be positively associated with sustainable success since these types of discussions may be more likely to incorporate diverse perspectives.
We collected all historical discussions between Wikipedians (i.e., discussers) on the talk page of each article and constructed two types of discussion features. The first type are linguistic features. We use pre-trained deep-learning models to measure those. Specifically, we measure the sentiment Hartmann et al. (2023), formality Babakov et al. (2023), politeness Danescu-Niculescu-Mizil et al. (2013); Wang and Jurgens (2018), toxicity Logacheva et al. (2022), and certainty Pei and Jurgens (2021) for each comment in a discussion. We then aggregate each, using the average, median, and percentage of high values. [^5] We use previously identified thresholds to define high values using each model To account for uneven contributions across the discussers of talk pages that might bias our measures, we also calculate the mean and median values for each discusser and then aggregate these values across all discussers.
The other type of features are related to the structure of the discussions. We list these 15 features in Appendix Table 3. Among those features are number-of-discussers, mean-responded-comments, and the discussers-gini Gini (1912), which is a measure of statistical dispersion commonly used to measure inequality. A Gini index of 0.0 indicates perfect equality where all editors contribute the same amount of content (i.e., comments in the case of discussions), while 1.0 indicates perfect inequality where a single individual generates all content.
@@
The way a team of Wikipedians engages with one another and with the edited article is crucial for the team’s success and resilience Ruprechter et al. (2020). We hypothesize that this is also correlated with sustainable success. This feature set, named $TC$, includes editors-gini ($TC_{1}$) and ip-based-edits-percentage ($TC_{2}$). $TC_{1}$ corresponds to the Gini inequality coefficient Gini (1912) among the Wikipedian editors. It is calculated in the same way as for discussers-gini (see previous paragraph) while taking into account article edit distribution among Wikipedians. $TC_{2}$ represents the fraction of revisions done by anonymous editors.